# Pre-Sorted Tsetlin Machine

### v1

@@ Abstract
This paper proposes a machine learning pre-sort stage to traditional supervised learning using Tsetlin Machines. Initially, N data-points are identified from the dataset using an expedited genetic algorithm to solve the maximum dispersion problem. These are then used as the initial placement to run the K-Medoid clustering algorithm. Finally, an expedited genetic algorithm is used to align N independent Tsetlin Machines by maximising hamming distance. For MNIST level classification problems, results demonstrate up to 10% improvement in accuracy, 383X reduction in training time and 86X reduction in inference time.
@@
Using an expedited genetic algorithm to ”conceptually arrange” the binarized datasets and identify N datapoints that provide the maximum dispersion based on hamming distance.
Using the K-Medoid clustering algorithm to sort same-class datapoints into N clusters based on hamming distance.
Using an expedited genetic algorithm to align the newly-formed clusters into N independent Tsetlin Machines.
@@ 3 Pre-Sort Architecture
The proposed pre-sort architecture can be split into three distinct stages: Binary Maximum Dispersion (M), Binary K-Medoid (K) and Genetic Class Alignment (G). These stages are run sequentially on the dataset. The result is a rapid sorting of data that may be split into N data subsets that are then distributed to smaller Tsetlin Machines. A high level overview of the pre-sort architecture may be seen in Figure 5.
@@ 3.1 Binary Maximum Dispersion
The purpose of this stage is to identify N datapoints within each class of the dataset that represent the maximum cumulative difference possible. This is commonly referred to as solving the maximum dispersion problem. For binarized datasets, there is usually no ordinality (one handwritten digit isn’t inherently greater or lesser than another). However, a ”distance” between two binarized datapoints can be measured using the hamming distance (the number of bits different between them). In this manner, the N maximally-distant datapoints may be identified and the remaining data ”conceptually arranged” by hamming distance between them. A visualiation of this process for N=8 datapoints may be seen in Figure 6.
To find the datapoints, an expedited genetic algorithm is used to determine the greatest cumulative hamming distance between N chosen datapoints. The final N datapoints need only be a rough estimate at this stage as they are updated by the next stage. The genetic algorithm therefore does not require execution until exhaustion and can be parallelized to further reduce run time.
@@ 3.2 Binary K-Medoid Algorithm
The Binary K-Medoid Algorithm performs the bulk of the work in the pre-sort architecture. Unlike the more common K-Means clustering algorithm, K-Medoid chooses an actual datapoint to cluster on. This makes the algorithm more robust when handling datasets with considerable outliers. Moreover, given that most binarized data has no ordinality, a calculated ”mean” would not provide any meaningful insights.
The algorithm first takes the N datapoints from the Binary Maximum Dispersion stage. These become the initial medoids. All non-medoid datapoints are then assigned to their nearest medoid based on hamming distance to form N clusters. Next, the cumulative distance from each datapoint is calculated to all other datapoints within the same cluster. The datapoint in each cluster with the smallest cumulative distance is then selected to be the new medoid for that cluster. All datapoints are then assigned to their nearest medoid and the process is iterated. The algorithm converges when the medoids are unchanged between iteration cycles.
@@ 3.3 Genetic Class Alignment
The dataset now consists of N clusters within each class. Given that the Tsetlin Machine functions by mining frequent sub-patterns and using these to differentiate between the classes, the optimal strategy is to align clusters based on maximal inter-class hamming distance and to distribute these to N separate Tsetlin Machines.
@@
To implement this concept, the final medoids from the K-Medoid stage (N medoids within each class) are placed with their class peers into 1D horizontal arrays. These arrays are then vertically stacked as shown in Figures 8 and 9 (such that each row represents one class). The total cumulative hamming distance between each medoid in a vertical stack is calculated and then summed for all vertical stacks to give a total hamming distance for the 2D configuration. An expedited genetic algorithm is used to search for the 2D configuration with the largest total hamming distance. The columns (containing one cluster from each class) are then used to separate the dataset in N smaller datasets. These are then processed by their own independent Tsetlin Machine.
@@ 4 Experiments
Datasets were classified with a varying number of medoids (2,4,8,16,32). Two runtime strategies were chosen. The first consisted of running both the baseline Tsetlin Machine algorithm and proposed pre-sort method for 100 epochs, then running inference. This determined the training time, inference time and accuracy. However, as the number of medoids was increased during the experiment, it became apparent that some pre-sort Tsetlin Machines reached 100% accuracy within 4-6 epochs, after which, it was superfluous to continue training. The second strategy was therefore to run an inference between each training cycle for these pre-sort Tsetlin Machines. When 100% accuracy was reached, the training was concluded. The training runtime for this method includes the inference time required to check the accuracy after each training epoch. The results presented show the fastest of these two methods for the pre-sort training time, as both are legitimate training strategies. All dataset/medoid points were performed 10 times using unique 80/20 train/test splits of the datasets and the average (mean) is presented. 4000 clauses per class were used for the baseline Testlin Machine implementation as suggested on the CAIR repository. For the pre-sorted Tsetlin Machines, 4000 clauses per class were used in total across all distributed Tsetlin Machines (eg N=2, 2000 clauses per class per machine). As the clustering is non-uniform, the number of clauses assigned to each distributed Tsetlin Machine was apportioned based on the total number of datapoints assigned to that Tsetlin Machine. Other options are available, but were not explored. All runs were performed single-threaded on a first-gen M1 Macbook Pro.
@@ 6 Discussion
It can be seen from Figure 11 that for all datasets, the training time is drastically reduced for the pre-sorted Tsetlin Machine with respect to the baseline implementation. There are several reasons for this. Principally, fewer clauses are run per training cycle. This is because in the baseline Tsetlin Machine, 4000 clauses are updated for every datapoint passed into a class. In pre-sorted Tsetlin Machine, only 1/N * 4000 clauses are updated on average per datapoint. This immediately results in a training runtime reduction by a factor of N. Another reason is that having pre-simplified the problem, the test accuracy reached 100% for some pre-sorted Tsetlin Machines after only a few epochs. It was therefore faster to adopt a strategy of testing for this condition and then terminating training rather than running for an extensive period of time.
It can be seen from Figure 12 that for all datasets, the inference time is drastically reduced for the pre-sorted Tsetlin Machine with respect to the baseline implementation. The reason for this is identical to the reduction in training time. The number of clauses required for an inference is reduced by a factor of N, lowering the computational complexity and therefore reducing the runtime.
As the entire pre-sort architecture may be parallelised, a further train/inference run time reduction of at least a factor of N is achievable against the results herein presented.
@@ 6.1 Momentary Mastery
As the original Tsetlin Machine is split out into smaller independent Tsetlin Machines in the proposed pre-sort scheme, a unique opportunity is presented for problems that only require ”one-shot” training and inference (eg training is performed, inference is performed, the results are stored and the model deleted). As each datapoint only requires one smaller independent Tsetlin Machine for training and inference and the datapoints are grouped together, only one of the smaller independent Tsetlin Machines needs to exist at any given time. This would result in a memory footprint reduction by a factor of N, provided that a single threaded training/inference speed was acceptable. This is a considerable saving for embedded/edge devices.
@@ 6.2 Hardware Impact
Whilst Section 3 provided sound conceptual reasoning behind why each of the three stages of the pre-sort architecture were chosen, perhaps more importantly, there is synergy with the hardware requirements of Tsetlin Machine. The heart of all three pre-sort stages is the hamming distance calculation: XOR followed by popcount. Whilst AND, XOR and XNOR (bitwise equality) are common logical operators, popcount is typically omitted in RISC architectures. Hence, extending these architectures with the additional popcount instruction improves the entire proposed pre-sort architecture, not just the Tsetlin Machine.
@@ 7 Final Comments
This work proposed and investigated a pre-sort architecture for Tsetlin Machines. Results were presented that demonstrated up to 10% improvement in accuracy, 383X reduction in training time and 86X reduction in inference time. A full delineation of how this was achieved was discussed, along with methods to reduce memory footprint and hardware considerations for improving performance.

### v2

@@ Abstract
This paper proposes a machine learning pre-sort stage to traditional supervised learning using Tsetlin Machines. Initially, K data-points are identified from the dataset using an expedited genetic algorithm to solve the maximum dispersion problem. These are then used as the initial placement to run the K-Medoid clustering algorithm. Finally, an expedited genetic algorithm is used to align K independent Tsetlin Machines by maximising hamming distance. For MNIST level classification problems, results demonstrate up to 10% improvement in accuracy, 383X reduction in training time and 99X reduction in inference time.
@@
Using an expedited genetic algorithm to ”conceptually arrange” the binarized datasets and identify K datapoints that provide the maximum dispersion based on hamming distance.
Using the K-Medoid clustering algorithm to sort same-class datapoints into K clusters based on hamming distance.
Using an expedited genetic algorithm to align the newly-formed clusters into K independent Tsetlin Machines.
@@ 3 Pre-Sort Architecture
The proposed pre-sort architecture can be split into three distinct stages: Binary Maximum Dispersion (M), Binary K-Medoid (K) and Genetic Class Alignment (G). These stages are run sequentially on the dataset. The result is a rapid sorting of data that may be split into K data subsets that are then distributed to smaller Tsetlin Machines. A high level overview of the pre-sort architecture may be seen in Figure 5.
@@ 3.1 Binary Maximum Dispersion
The purpose of this stage is to identify K datapoints within each class of the dataset that represent the maximum cumulative difference possible. This is commonly referred to as solving the maximum dispersion problem. For binarized datasets, there is usually no ordinality (one handwritten digit isn’t inherently greater or lesser than another). However, a ”distance” between two binarized datapoints can be measured using the hamming distance (the number of bits different between them). In this manner, the K maximally-distant datapoints may be identified and the remaining data ”conceptually arranged” by hamming distance between them. A visualiation of this process for K=8 datapoints may be seen in Figure 6.
To find the datapoints, an expedited genetic algorithm is used to determine the greatest cumulative hamming distance between K chosen datapoints. The final K datapoints need only be a rough estimate at this stage as they are updated by the next stage. The genetic algorithm therefore does not require execution until exhaustion and can be parallelized to further reduce run time.
@@ 3.2 Binary K-Medoid Algorithm
The Binary K-Medoid Algorithm performs the bulk of the work in the pre-sort architecture. Unlike the more common K-Means clustering algorithm, K-Medoid chooses an actual datapoint to cluster on. This makes the algorithm more robust when handling datasets with considerable outliers. Moreover, given that most binarized data has no ordinality, a calculated ”mean” would not provide any meaningful insight.
The algorithm first takes the K datapoints from the Binary Maximum Dispersion stage. These become the initial medoids. All non-medoid datapoints are then assigned to their nearest medoid based on hamming distance to form K clusters. Next, the cumulative distance from each datapoint is calculated to all other datapoints within the same cluster. The datapoint in each cluster with the smallest cumulative distance is then selected to be the new medoid for that cluster. All datapoints are then assigned to their nearest medoid and the process is iterated. The algorithm converges when the medoids are unchanged between iteration cycles.
@@ 3.3 Genetic Class Alignment
The dataset now consists of K clusters within each class. Given that the Tsetlin Machine functions by mining frequent sub-patterns and using these to differentiate between the classes, the optimal strategy is to align clusters based on maximal inter-class hamming distance and to distribute these to K separate Tsetlin Machines.
@@
To implement this concept, the final medoids from the K-Medoid stage (K medoids within each class) are placed with their class peers into 1D horizontal arrays. These arrays are then vertically stacked as shown in Figures 8 and 9 (such that each row represents one class). The total cumulative hamming distance between each medoid in a vertical stack is calculated and then summed for all vertical stacks to give a total hamming distance for the 2D configuration. An expedited genetic algorithm is used to search for the 2D configuration with the largest total hamming distance. The columns (containing one cluster from each class) are then used to separate the dataset in K smaller datasets. These are then processed by their own independent Tsetlin Machine.
@@ 4 Experiments
Datasets were classified with a varying number of medoids (2,4,8,16,32). Two runtime strategies were chosen. The first consisted of running both the baseline Tsetlin Machine algorithm and proposed pre-sort method for 100 epochs, then running inference. This determined the training time, inference time and accuracy. However, as the number of medoids was increased during the experiment, it became apparent that some pre-sort Tsetlin Machines reached 100% accuracy within 4-6 epochs, after which, it was superfluous to continue training. The second strategy was therefore to run an inference between each training cycle for these pre-sort Tsetlin Machines. When 100% accuracy was reached, the training was concluded. The training runtime for this method includes the inference time required to check the accuracy after each training epoch. The results presented show the fastest of these two methods for the pre-sort training time, as both are legitimate training strategies. All dataset/medoid points were performed 10 times using unique 80/20 train/test splits of the datasets and the average (mean) is presented. 4000 clauses per class were used for the baseline Testlin Machine implementation as suggested on the CAIR repository. For the pre-sorted Tsetlin Machines, 4000 clauses per class were used in total across all distributed Tsetlin Machines (eg K=2, 2000 clauses per class per machine). As the clustering is non-uniform, the number of clauses assigned to each distributed Tsetlin Machine was apportioned based on the total number of datapoints assigned to that Tsetlin Machine. Other options are available, but were not explored. All runs were performed single-threaded on a first-gen M1 Macbook Pro.
@@ 6 Discussion
It can be seen from Figure 11 that for all datasets, the training time is drastically reduced for the pre-sorted Tsetlin Machine with respect to the baseline implementation. There are several reasons for this. Principally, fewer clauses are run per training cycle. This is because in the baseline Tsetlin Machine, 4000 clauses are updated for every datapoint passed into a class. In pre-sorted Tsetlin Machine, only 1/K * 4000 clauses are updated on average per datapoint. This immediately results in a training runtime reduction by a factor of K. Another reason is that having pre-simplified the problem, the test accuracy reached 100% for some pre-sorted Tsetlin Machines after only a few epochs. It was therefore faster to adopt a strategy of testing for this condition and then terminating training rather than running for an extensive period of time.
It can be seen from Figure 12 that for all datasets, the inference time is drastically reduced for the pre-sorted Tsetlin Machine with respect to the baseline implementation. The reason for this is identical to the reduction in training time. The number of clauses required for an inference is reduced by a factor of K, lowering the computational complexity and therefore reducing the runtime.
As the entire pre-sort architecture may be parallelised, a further train/inference run time reduction of at least a factor of K is achievable against the results herein presented simply by running multi-threaded.
@@ 6.1 Momentary Mastery
As the original Tsetlin Machine is split out into smaller independent Tsetlin Machines in the proposed pre-sort scheme, a unique opportunity is presented for problems that only require ”one-shot” training and inference (eg training is performed, inference is performed, the results are stored and the model deleted). As each datapoint only requires one smaller independent Tsetlin Machine for training and inference and the datapoints are grouped together, only one of the smaller independent Tsetlin Machines needs to exist at any given time. This would result in a memory footprint reduction by a factor of K, provided that a single threaded training/inference speed was acceptable. This is a considerable saving for embedded/edge devices.
@@ 6.2 Hardware Impact
Whilst Section 3 provided sound conceptual reasoning behind why each of the three stages of the pre-sort architecture were chosen, perhaps more importantly, there is synergy with the hardware requirements of Tsetlin Machine. The heart of all three pre-sort stages is the hamming distance calculation: XOR followed by popcount. Whilst AND, XOR (bitwise inequality) and XNOR (bitwise equality) are common logical operators, popcount is typically omitted in RISC architectures. Hence, extending these architectures with the additional popcount instruction improves the entire proposed pre-sort architecture, not just the Tsetlin Machine.
@@ 7 Final Comments
This work proposed and investigated a pre-sort architecture for Tsetlin Machines. Results were presented that demonstrated up to 10% improvement in accuracy, 383X reduction in training time and 99X reduction in inference time. A full delineation of how this was achieved was discussed, along with methods to reduce memory footprint and hardware considerations for improving performance.